OpenStudy (anonymous):

is the inverse sine of sine of (-(4pi)/9)) equal pi/9?

OpenStudy (anonymous):

sin(-x) = -sinx

OpenStudy (anonymous):

-sin^-1 (sin 4pie/9)

OpenStudy (anonymous):

so -4pie/9

OpenStudy (anonymous):

Why not just -pi/9?

OpenStudy (anonymous):

But pi/9 is in that range?

myininaya (myininaya):

Oh. Wait It is.

OpenStudy (anonymous):

What is?

myininaya (myininaya):

lol. So yeah \[\frac{-4 \pi }{9} \in [\frac{-\pi}{2}, \frac{\pi}{2}] =>\sin^{-1}(\sin(\frac{-4 \pi}{9}))=\frac{-4 \pi}{9} \]

myininaya (myininaya):

\[-.444 \pi \in [-.5 \pi , .5 \pi]\] Since \[-.444 \in [-.5,.5]\]

myininaya (myininaya):

-4/9 is approximately -.444 :)